OpenStudy (anonymous):

Please help...I have been doing homework for 4+ hours, and my brain is fried...Can someone do the steps and show me how to do this? Award At End. Given r(t), find the following: r(t) = ... Write a(t) in terms of its tangential and normal components.

ganeshie8 (ganeshie8):

divide that by its magnitude to get the unit tangent vector

OpenStudy (anonymous):

Divide what? <2t, tsin(t), tcos(t)>

OpenStudy (anonymous):

How do I solve for the magnitude ?

ganeshie8 (ganeshie8):

\[\large \hat{T(t)} = \dfrac{1}{t\sqrt{5}}\langle 2t, ~t\sin t, ~t\cos t\rangle\]

ganeshie8 (ganeshie8):

magnitude = \(\large \sqrt{(2t)^2 + (t\sin t)^2 + (t\cos t)^2} = t\sqrt{5}\)

ganeshie8 (ganeshie8):

Oh well, u can cancel \(t\) as well

ganeshie8 (ganeshie8):

Unit tangent vector : \[\large \hat{T(t)} = \dfrac{1}{\sqrt{5}}\langle 2, ~\sin t, ~\cos t\rangle\]

OpenStudy (anonymous):

So its < 2/Sqr(5) , sin(t)/Squr(5) , cos(5)/Sqr(5) >

ganeshie8 (ganeshie8):

yes but thats just the unit tangent vector, not the component of acceleration

OpenStudy (anonymous):

Haha yeah I saw that extra t and canceled it out. :)

OpenStudy (anonymous):

Okay and to get acceleration whats the formula?

OpenStudy (anonymous):

I have it written down somewhere in my notes...Gotta rummage through them /.\

ganeshie8 (ganeshie8):

velocity = first derivative acceleration = second derivative Simple.

OpenStudy (anonymous):

So basically we have to solve for velocity first. So we need to take the derivative of r(t) = <t^2, sin(t) - tcos(t), cos(t) + tsin(t)>

ganeshie8 (ganeshie8):

yes

OpenStudy (anonymous):

Then we take the derivative of that! Hah!

OpenStudy (anonymous):

Okay hold on...Let me try this..

OpenStudy (anonymous):

So < 2 , sin(t) + tcos(t) , cos(t) - tsin(t) > ??

OpenStudy (anonymous):

@ganeshie8 Do I just remove the 2 because this is the second derivative?

ganeshie8 (ganeshie8):

nope, thats the acceleration vector.

ganeshie8 (ganeshie8):

next, find the tangent and normal components

ganeshie8 (ganeshie8):

for tangent component, do below : 1) find the magnitude of acceleration 2) multiply it by the unit tangent vector

ganeshie8 (ganeshie8):

\[\overrightarrow{A(t)} = \langle 2 , \sin(t) + t\cos(t) , \cos(t) - t\sin(t) \rangle \] \[|\overrightarrow{A(t)} | = ?\]

OpenStudy (anonymous):

Wait I though..Magnitude of acceleration = Change of velocity / Time interval

OpenStudy (anonymous):

thought*

ganeshie8 (ganeshie8):

and ur definition of Magnitude of velocity is Change of position / Time interval ?

ganeshie8 (ganeshie8):

what u have there are average acceleration and average velocity - for instantaneous acceleration u need to use derivatives.

ganeshie8 (ganeshie8):

\[\overrightarrow{A(t)} = \langle 2 , \sin(t) + t\cos(t) , \cos(t) - t\sin(t) \rangle\] can u find its magnitude ?

OpenStudy (anonymous):

Ohhh okay...I'm kinda lost now. :/

ganeshie8 (ganeshie8):

well, you haven't answered my question on whether u can find the magnitude of given Acceleration vector :/

OpenStudy (anonymous):

Yeah I am trying to...Im getting lost...:/

OpenStudy (anonymous):

To find the magnitude of A(t) what do we do? What formula?

ganeshie8 (ganeshie8):

magnitude of <x, y, z> = sqrt(x^2 + y^2 + z^2)

OpenStudy (anonymous):

So how would I find the magnitude of A(t). Do I just square < 2 , sin(t) + tcos(t) , cos(t) - tsin(t) > ?

OpenStudy (anonymous):

After I times x,y, and z by ^2 of course. @ganeshie8

ganeshie8 (ganeshie8):

\(\large \overrightarrow{A(t)} = \langle \color{green}{2} , \color{blue}{\sin(t) + t\cos(t)} , \color{red}{\cos(t) - t\sin(t)} \rangle \) Magnitude = \(\large \sqrt{( \color{green}{2})^2 + ( \color{blue }{\sin(t)+t\cos(t)})^2 + ( \color{red}{\cos(t) - t\sin(t)})^2}\)

ganeshie8 (ganeshie8):

simplify^

OpenStudy (anonymous):

Uhhh.... 4 + t^2 cos^2(t)+2 t cos(t) sin(t)+sin^2(t) + cos^2(t)-2 t cos(t) sin(t)+t^2 sin^2(t)

OpenStudy (anonymous):

I can simplify that more I think..

OpenStudy (anonymous):

Ohh...I expanded it out.

ganeshie8 (ganeshie8):

So, the tangent component of \(\large \overrightarrow{A(t)} \) is : \(|\large \overrightarrow{A(t)}| *\hat{T(t)} = \sqrt{t^2+5} * \dfrac{1}{\sqrt{5}}\langle 2, ~\sin t, ~\cos t\rangle \)

ganeshie8 (ganeshie8):

you may find the normal component of \(\large \overrightarrow{A(t)}\) same way : u just need to find the unit normal vector and multiply it by the magnitude of \(\large \overrightarrow{A(t)}\) good luck :)